Two samples of a monatomic ideal gas are in separate containers at the same conditions of pressure, volume, and temperature ( help please!!!
Answer #1

Sample 1: (P1,V1) → (P1, V2) → (P2, V2) Work done (w) = w1 + w2 = - Pext(V+ + V;) + 0 = - 1.00atm(2.00 - 1.00)L = -1.00atm-L